calculate the number of moles of sodium oxide that should be produced when 26.50 grams of sodium metal reacts with oxygen gas. show work!


1
Expert's answer
2021-04-14T05:55:25-0400

Solution:

Balanced chemical equation:

4Na(s) + O2(g) → 2Na2O(s)

According to the equation above: n(Na)/2 = n(Na2O)


Moles of Na = Mass of Na / Molar mass of Na

The molar mass of Na is 23 g mol-1.

Hence,

Moles of Na = 26.50 g / 23 g mol-1= 1.152 mol

n(Na) = 1.152 mol


n(Na2O) = n(Na)/2 = 1.152 / 2 = 0.576 mol

Moles of Na2O = 0.576 mol


OR (short form solution):

(26.50 g Na) × (1 mol Na / 23 g Na) × (2 mol Na2O / 4 mol Na) = 0.576 mol Na2O


Answer: 0.576 moles of sodium oxide (Na2O) should be produced.